### v1

@@ 1 Introduction
Advancements in brain imaging technologies, especially Functional Magnetic Resonance Imaging (fMRI) of Blood-Oxygen-Level-Dependent (BOLD) signals, have enabled precise and quantitative measures of brain activity, leading to a significant boost in neuroscience research [2]. Among various brain modeling approaches, causation models have recently received a revamped interest [3, 4] in recent years, due to their capability to reveal complex relationships without multi-scale brain networks. These models have been extended in multiple directions to understand the brain’s underlying processes [5], explaining how these processes support cognition and behavior [6, 7, 8], and propose new methods to regulate brain activity [9, 10] for enhancing normal functions and mitigating disorders [11].
Inspired by these developments, this paper explores the use of causal dynamic models to extract informative features from the brain for fMRI fingerprinting. Specifically, we aim to leverage causal signatures derived from fMRI time-series data to identify both the individual subject (subject fingerprinting) and the cognitive task being performed (task fingerprinting). A major challenge in this process arises from the variability across subjects and tasks, along with the limited number of recordings available for each subject-task combination [12]. Our approach focuses on uncovering and utilizing latent causal signatures in the brain. While this requires more specialized model structures, it can offer improved data efficiency. This distinguishes our method from traditional correlation-based approaches [13, 14, 10, 7, 15, 16, 17, 18, 19], which are limited in their ability to capture the directional, evolving cause-and-effect relationships that are fundamental to cognitive dynamics.
@@ 2 Related work
In neuroscience research, one strategy to address this challenge is to use the functional connectome (FC) [12]. The FC measures the correlations of the fMRI time-series data among pair-wise brain regions during rest or task conditions, and it has been successfully utilized as individual-specific neural signatures to classify various cognitive states and neurological disorders[24, 18]. Another common brain-network modeling technique is Independent Component Analysis (ICA). This approach separates fMRI data into spatially distinct, statistically independent components corresponding to different brain networks [25, 26]. It has proven especially useful for revealing resting-state networks and explaining their roles in cognitive function and pathology [27, 28, 29]. Although FC and ICA have achieved notable success in many applications, they still have several limitations. In particular, they do not fully capture the dynamic nature of brain activity, as they lack the temporal resolution needed to detail interactions among regions, and cannot represent the directionality of connections between regions [24, 30, 14, 16]. Each of these aspects, dynamics, temporal detail, and directionality, is essential for understanding the cause-and-effect interactions within brain networks.
@@
In contrast to static correlational methods, causality-based approaches such as Dynamic Causal Modeling (DCM) and Granger causality offer alternative insights into patterns of brain interactions. However, these directed connectivity methods have been relatively under-explored in the context of fMRI fingerprinting. Analyzing causality in fMRI entails examining directional influences among brain regions over time, which provides a window into the brain’s dynamic modes of interaction. For example, DCM uses a Bayesian framework to model interactions among latent neural states with bilinear differential equations, which describes how activity in one region influences others over time [31, 32]. Similarly, Granger causality evaluates whether one time series can predict another, thereby identifying potential directed relationships between neural signals [33, 34, 35]. To further enhance these causal models and better capture complex neural dynamics, recent research has introduced several modifications. For instance, threshold-based structures have been incorporated to reflect the saturation of neural activation [36, 37, 38], and multifactorial dynamics have been proposed to represent systems that operate across multiple time scales [39].
Beyond these correlation-based or causality-based methodologies, classical machine‑learning algorithms also remain popular. Random Forest (RF) [40] and Support Vector Machines (SVM) [41] provide high interpretability and easy deployment across diverse datasets. Recent studies on fMRI fingerprinting have also explored a variety of advanced techniques to capture different facets of brain dynamics and connectivity patterns. These include deep‑learning architectures. Deep Neural Networks (DNN) capture high‑dimensional connectivity patterns [42]. Deep Bidirectional Recurrent Neural Networks (DBRNN) model temporal dynamics in brain signals [43]. Brain Attend and Decode (BAnD) applies attention mechanisms to highlight salient brain regions [44].
@@ 3.1 A Two-timescale State-space Model for Causal Dynamics
Figure 1: Two–timescale partition of a single fMRI sampling period $[(k-1)\Delta t,\,k\Delta t]$. The interval is split into a fast sub‑interval of duration $\tau$ (short solid red) and a slow sub‑interval of duration $\Delta t-\tau$ (long solid red), with $0<\tau\ll\Delta t$. Driven by the slower input $u(k-1)$ and the current input $u(k)$, the state evolves from the previous scan $x(k-1)$, passes the intermediate point $z(k)=x(k-\tau)$, and reaches the current scan $x(k)$.
@@
Here, $Q\in\mathbb{R}^{m\times m}$ captures the fast sub-interval (concurrent) interaction among the neuronal states; $A\in\mathbb{R}^{m\times m}$ encodes the slow sub-interval (cross-lagged) transition from $x(k-1)$; $B_{1}\,u(k)$ and $B_{2}\,u(k-1)$ incorporate the inputs into the system for fast and slow dynamics, respectively.
@@
The slow dynamics are approximated using an explicit Euler approximation from $x(k-1)$ to an intermediate state $z(k)$:
@@
where $\lambda>0$ controls the strength of the regularization to prevent overfitting, and $\|\cdot\|_{F}$ denotes the Frobenius norm [47] to minimize the residue in (3.1).
By solving (3.1), we transform a complex fMRI time-series data $d(t)=\{x(t),u(t)\}$ into a structured and meaningful representation ${R}=[Q~{}A~{}B_{1}~{}B_{2}]\in\mathbb{R}^{m\times(2m+2n)}$ of brain dynamics signature, as visualized in Fig. 2 . This transformation not only makes the data more accessible for subsequent analysis but also reveals the patterns of ‘spatio-temporal’ causality relationships that underpin cognitive processes. In the following, we will build on $R$ to perform subject and task fingerprints.
@@ 3.2 Subject Causal Fingerprint via Modal Decomposition and Projection
^$\widehat{\mathcal{R}}_{h}=\{R_{s,h}^{D}\},\quad s\in\mathcal{S}.$
where the superscript $D$ denotes data samples with known subject identities. We assume that the dataset includes at least two recordings for every subject-task pair. This allows one sample to be used in the reference set ^$\widehat{\mathcal{R}}_{h}$, while the remaining samples serve as query instances for evaluation. Given a query sample with causal signature $R_{h}^{Q}$ of an unknown identity, the objective is to identify the subject label $s$ by comparing it against the labeled reference set $R_{s,h}^{D}$.
The key challenge for subject causal fingerprinting lies in the limited data for each subject, as the fMRI data collection is time-consuming. Essentially, we need to solve a one-shot classification problem [23], i.e., only one sample per subject is available to determine their identity from a large number of candidates. To address this, we leverage the control-theoretic interpretation of the model (1), in which the causal system signature $R=[Q~{}A~{}B_{1}~{}B_{2}]$ naturally encodes the system’s dynamic modes [48, Chapter 12], reflecting invariant subspaces of brain dynamics. These dynamic modes serve as augmented features for robust subject identification.
@@
[…]
where $\mathcal{T}$ is a feature set composing $[\widehat{T},~{}\bar{T}]$, or a subset of these features, depending on which features best capture inter-subject variation. A detailed discussion on this will be given in the Results section. The distance metric $\text{Dist}(\cdot,\cdot)$ is computed as the permutation-aligned pairwise vector similarity (Euclidean dot product) through the projection between the columns of the transformation matrices. Note that the eigenvalue matrices $\bar{\Lambda}$ and $\widehat{\Lambda}$ quantify the significance of each mode in the dynamics. We, however, do not incorporate them directly in the fingerprinting task. Instead, our method treats all dynamic modes as equally informative, regardless of their eigenvalues. This contrasts with dimensionality-reduction methods like Principal Component Analysis (PCA), which retain only the most significant components. Our method, via (11), operates in the full space spanned by all dynamic modes, under the assumption that when different subjects perform the same fMRI task, their major dynamic modes might be similar but vary in certain minor dynamic modes.
@@ 3.3 Task Causal Fingerprint via Graph Neural Network
^=$\widehat{\mathcal{R}}=\{R^{D}_{s,h}\},\quad s\in\mathcal{S},\text{ and }h\in\mathcal{H}$
@@
Building on this input structure, we employ a five-layer Graph Neural Network architecture tailored for task fingerprinting from fMRI data. The network starts with a GATv2Conv layer that re-weights each node’s neighbours, emphasising the connections most relevant to the current task [51]. A parallel linear skip path is summed with this output so the original signal remains visible to deeper layers, which improves gradient flow and protects against over-smoothing [52]. TopKPooling then selectively retains the most informative nodes. It reduces graph complexity and directs computational resources toward the most relevant substructures. This sparsification step improves both efficiency and performance. A second GATv2Conv block, again accompanied by a skip connection, further refines the retained features and maps them into a more discriminative space for task recognition. Global mean pooling aggregates these refined node features across the graph, generating a single, comprehensive graph-level representation vector. A final linear head followed by log-softmax converts this vector into task probabilities for training and evaluation. The overall model structure is visualized in Figure 3.
The model is trained using supervised learning with data batches drawn from $\widehat{\mathcal{R}}$. The rows of $R^{D}_{s,h}$ serve as the input features for each node, while $h$ is used as the corresponding fMRI task label. Detailed training parameters are provided in the Appendix.
@@ 4 Results
Method CM+MD&P CM+CoR CM+FN CM+GNN FC+CoR FC+MD&P FC+FN FC+GNN
Rest1 LR 95.396% 58.397% 42.251% $<5\%$ 45.439% 69.736% 33.952% $<5\%$
Rest2 LR 96.334% 61.295% 47.175% $<5\%$ 45.436% 70.929% 37.574% $<5\%$
Rest1 RL 94.970% 57.118% 41.333% $<5\%$ 41.603% 67.945% 38.698% $<5\%$
Rest2 RL 94.203% 58.978% 43.463% $<5\%$ 46.803% 67.775% 36.737% $<5\%$
@@ 4.1 Dataset and Modeling
To ensure the quality and consistency of the data, we performed a series of preprocessing steps. First, global signals were regressed out from each time series based on insights from [56]. Next, we smoothed the time series using a Savitzky-Golay filter with a window size of 11, fitting a local polynomial to the data for enhanced temporal consistency. Finally, z-score standardization was applied to normalize the data.
Figure 6: Left: Subject Fingerprinting based on CM+MD&P: The identification accuracy is compared with (black line) using the FC+CoR method. Right: Task Fingerprinting based on CM+GNN: The figure shows the correct/incorrect classification for the same/different fMRI tasks.

### v2

@@ 1 Introduction
Advancements in brain imaging technologies, especially Functional Magnetic Resonance Imaging (fMRI) of Blood-Oxygen-Level-Dependent (BOLD) signals, have enabled precise and quantitative measures of brain activity, leading to a significant boost in neuroscience research [2]. Among various brain modeling approaches, causation models have recently received a revamped interest [3, 4] in recent years, due to their capability to reveal complex relationships without multi-scale brain networks. These models have been extended in multiple directions to understand the brain’s underlying processes [5], explaining how these processes support cognition and behavior [6, 7, 8], and propose new methods to regulate brain activity [9, 10] for enhancing normal functions and mitigating disorders [11]. However, most fingerprinting pipelines still rely on static, symmetric connectivity maps and thus miss the directed, multi-timescale dynamics that underlie neural communication.
Inspired by these developments, this paper explores the use of causal dynamic models to extract informative features from the brain for fMRI fingerprinting. Specifically, we aim to leverage causal signatures derived from fMRI time-series data to identify both the individual subject (subject fingerprinting) and the cognitive task being performed (task fingerprinting). A major challenge in this process arises from the variability across subjects and tasks, along with the limited number of recordings available for each subject-task combination [12]. In other words, the problem combines heterogeneity (across subjects/tasks) with data sparsity, which calls for models that are both interpretable and sample-efficient. Our approach focuses on uncovering and utilizing latent causal signatures in the brain. While this requires more specialized model structures, it can offer improved data efficiency. This distinguishes our method from traditional correlation-based approaches [13, 14, 10, 7, 15, 16, 17, 18, 19], which are limited in their ability to capture the directional, evolving cause-and-effect relationships that are fundamental to cognitive dynamics.
@@ 2.1 Static connectivity models
In neuroscience research, one strategy to address this challenge is to use the functional connectome (FC) [12]. The FC measures the correlations of the fMRI time-series data among pair-wise brain regions during rest or task conditions, and it has been successfully utilized as individual-specific neural signatures to classify various cognitive states and neurological disorders [24, 18]. However, FC is static and undirected, thus missing causal directionality. Another common brain-network modeling technique is Independent Component Analysis (ICA). This approach separates fMRI data into spatially distinct, statistically independent components corresponding to different brain networks [25, 26]. It has proven especially useful for revealing resting-state networks and explaining their roles in cognitive function and pathology [27, 28, 29]. Although FC and ICA have achieved notable success in many applications, they still have several limitations. In particular, they do not fully capture the dynamic nature of brain activity, as they lack the temporal resolution needed to detail interactions among regions, and cannot represent the directionality of connections between regions [24, 30, 14, 16]. Each of these aspects, dynamics, temporal detail, and directionality, is essential for understanding the cause-and-effect interactions within brain networks.
2.2 Causal connectivity models
@@
In contrast to static correlational methods, causality-based approaches such as Dynamic Causal Modeling (DCM) and Granger causality offer alternative insights into patterns of brain interactions. However, these directed connectivity methods have been relatively under-explored in the context of fMRI fingerprinting. Analyzing causality in fMRI entails examining directional influences among brain regions over time, which provides a window into the brain’s dynamic modes of interaction. For example, DCM uses a Bayesian framework to model interactions among latent neural states with bilinear differential equations, which describes how activity in one region influences others over time [31, 32]. Similarly, Granger causality evaluates whether one time series can predict another, thereby identifying potential directed relationships between neural signals [33, 34, 35]. Yet such directed methods remain under-explored for fMRI fingerprinting. Most Dynamic Causal Modeling (DCM) or Granger-causality studies restrict the analysis to a small number of regions of interest (ROIs) and assume multiple sessions per subject. Consequently, it is unclear how well these approaches scale to whole-cortex parcellations with few runs per subject, or to joint subject and task fingerprinting. To further enhance these causal models and better capture complex neural dynamics, recent research has introduced several modifications. For instance, threshold-based structures have been incorporated to reflect the saturation of neural activation [36, 37, 38], and multifactorial dynamics have been proposed to represent systems that operate across multiple time scales [39].
2.3 Learning-based approaches
Beyond these correlation-based or causality-based methodologies, classical machine-learning algorithms also remain popular. Random Forest (RF) [40] and Support Vector Machines (SVM) [41] provide high interpretability and easy deployment across diverse datasets. Recent studies on fMRI fingerprinting have also explored a variety of advanced techniques to capture different facets of brain dynamics and connectivity patterns. These include deep-learning architectures. Deep Neural Networks (DNN) capture high-dimensional connectivity patterns [42]. Deep Bidirectional Recurrent Neural Networks (DBRNN) model temporal dynamics in brain signals [43]. Brain Attend and Decode (BAnD) applies attention mechanisms to highlight salient brain regions [44]. A recent variant, TCN–BiLSTM, combines dilated Temporal Convolutional Networks (TCNs) with bidirectional Long Short-Term Memory (BiLSTM) layers to model short- and long-range dependencies in fMRI sequences [45]. However, it still operates on undirected regional time series and lacks an explicit multi-timescale causal framework, which limits its ability to characterize directed network interactions. Geometry- and topology-based pipelines have also been proposed. Tangent-Space Fingerprinting (TSF) maps functional-connectivity matrices to the Riemannian tangent space of the symmetric positive-definite (SPD) manifold, capturing global correlation geometry while still treating connectivity as static and undirected [46]. The PH-H0 landscape describes networks via zero-dimensional persistent homology (H0), providing a noise-robust topological summary yet discarding temporal scale and edge direction [47]. Classical classifiers such as Random Forests and Support Vector Machines are trained on pre-specified connectivity vectors. Without directionality or temporal context, their accuracy soon plateaus and lags behind causal models [40, 41]. End-to-end deep networks (DNN, DBRNN, BAnD) can harvest richer patterns, yet they require large training sets and remain hard to interpret. Geometry-based TSF and topology-based PH-H0 offer elegant dimensionality reduction, but both treat connectivity as static and undirected [47]. Across these lines of work, directed interactions and the multiple time scales that shape cognition are largely ignored. Our method fills this gap: a single scan yields compact causal signatures that encode directionality and tempo, and a lightweight GNN turns them into accurate fingerprints while preserving biophysical meaning.
@@ 3.1 A Two-timescale State-space Model for Causal Dynamics
Figure 1: Two–timescale partition of a single fMRI sampling period $(k-1)\Delta t,\,k\Delta t]$. The interval is split into a fast sub‑interval of duration $\tau$ (short solid red) and a slow sub‑interval of duration $\Delta t-\tau$ (long solid red), with $0<\tau\ll\Delta t$. Driven by the slower input u(k1)u(k-1) and the current input $u(k)$, the state evolves from the previous scan x(k1)x(k-1), passes the intermediate point $z(k)=x(k-\tau)$, and reaches the current scan $x(k)$.
@@
Here, $Q\in\mathbb{R}^{m\times m}$ captures the fast sub-interval (concurrent) interaction among the neuronal states; $A\in\mathbb{R}^{m\times m}$ encodes the slow sub-interval (cross-lagged) transition from x(k1)x(k-1); $B_{1}\,u(k)$ and $B_{2}\,u(k-1)$ incorporate the inputs into the system for fast and slow dynamics, respectively.
@@
The slow dynamics are approximated using an explicit Euler approximation from x(k1)x(k-1) to an intermediate state $z(k)$:
@@
Here $Q,A\in\mathbb{R}^{m\times m}$, $B_{1},B_{2}\in\mathbb{R}^{m\times n}$, $X^{1:T},X^{0:T-1}\in\mathbb{R}^{m\times T}$, and $U^{0:T-1},U^{1:T}\in\mathbb{R}^{n\times T}$, where $\lambda>0$ controls the strength of the regularization to prevent over-fitting, and $\|\cdot\|_{F}$ denotes the Frobenius norm (for any matrix $M\!\in\!\mathbb{R}^{p\times q}$, $\|M\|_{F}=\sqrt{\sum_{i=1}^{p}\sum_{j=1}^{q}M_{ij}^{2}}$[50]. The residual term $\bigl\|(Q-I)X^{1:T}+AX^{0:T-1}+B_{1}U^{0:T-1}+B_{2}U^{1:T}\bigr\|_{F}$ is the Frobenius norm of an $m\times T$ matrix, which aims to extract system parameters as causal signatures by minimizing the difference between the system’s measured next time state with the next time state predicted by our two-time scale model. The $\|Q\|_{F}$, $\|A\|_{F}$, $\|B_{1}\|_{F}$ and $\|B_{2}\|_{F}$ regularization terms to prevent over-fitting. By solving (3.1), we transform a complex fMRI time-series data $d(t)=\{x(t),u(t)\}$ into a structured and meaningful representation ${R}=[Q~A~B_{1}~B_{2}]\in\mathbb{R}^{m\times(2m+2n)}$ of brain dynamics signature, as visualized in Fig. 2 . This transformation not only makes the data more accessible for subsequent analysis but also reveals the patterns of ‘spatio-temporal’ causality relationships that underpin cognitive processes. In the following, we will build on $R$ to perform subject and task fingerprints.
@@ 3.2 Subject Causal Fingerprint via Modal Decomposition and Projection
$\widehat{\mathcal{R}}_{h}=\{R_{s,h}^{D}\},\quad s\in\mathcal{S}.$
where the superscript $D$ denotes data samples with known subject identities. We assume that the dataset includes at least two recordings for every subject-task pair. This allows one sample to be used in the reference set $\widehat{\mathcal{R}}_{h}$, while the remaining samples serve as query instances for evaluation. Given a query sample with causal signature $R_{h}^{Q}$ of an unknown identity, the objective is to identify the subject label $s$ by comparing it against the labeled reference set $R_{s,h}^{D}$.
The key challenge for subject causal fingerprinting lies in the limited data for each subject, as the fMRI data collection is time-consuming. Essentially, we need to solve a one-shot classification problem [23], i.e., only one sample per subject is available to determine their identity from a large number of candidates. To address this, we leverage the control-theoretic interpretation of the model (1), in which the causal system signature $R=[Q~A~B_{1}~B_{2}]$ naturally encodes the system’s dynamic modes [51, Chapter 12], reflecting invariant subspaces of brain dynamics. These dynamic modes serve as augmented features for robust subject identification.
@@
$\underset{s}{\arg\min}\quad\text{Dist}(\mathcal{T}_{R_{h}^{Q}},~\mathcal{T}_{R_{s,h}^{D}})$ (11)
where $\mathcal{T}$ is a feature set composing $\widehat{T},~\bar{T}]\in\mathbb{R}^{m\times 2m}$, or a subset of these features, depending on which features best capture inter-subject variation. A detailed discussion on this will be given in the Results section. The distance metric $\text{Dist}(\cdot,\cdot)$ is computed as the permutation-aligned pairwise vector distance, which compares two sets of vectors, where the elements are unordered. The key idea is to compute the cosine distance between elements in the two sets after finding an optimal permutation alignment that minimize overall distances, specifically, if defining $\mathcal{T}_{R_{h}^{Q}}=[x_{1},\cdots,x_{2m}],~\mathcal{T}_{R_{s,h}^{D}}=[y_{1},\cdots,y_{2m}]$,
$\text{Dist}(\mathcal{T}_{R_{h}^{Q}},~\mathcal{T}_{R_{s,h}^{D}})=\min_{\pi\in\mathcal{Z}}\sum_{i=1}^{n}\left(1-\frac{x_{i}^{\top}y_{\pi(i)}}{\|x_{i}\|\|y_{\pi(i)}\|}\right)$
where $\mathcal{Z}$ denotes the set of all permutations of $\{1,2,\ldots,2m\}$. Note that the eigenvalue matrices $\bar{\Lambda}$ and $\widehat{\Lambda}$ quantify the significance of each mode in the dynamics. We, however, do not incorporate them directly in the fingerprinting task. Instead, our method treats all dynamic modes as equally informative, regardless of their eigenvalues. This contrasts with dimensionality-reduction methods like Principal Component Analysis (PCA), which retain only the most significant components. Our method, via (11), operates in the full space spanned by all dynamic modes, under the assumption that when different subjects perform the same fMRI task, their major dynamic modes might be similar but vary in certain minor dynamic modes.
@@ 3.3 Task Causal Fingerprint via Graph Neural Network
$\widehat{\mathcal{R}}=\{R^{D}_{s,h}\},\quad s\in\mathcal{S},\text{ and }h\in\mathcal{H}$
@@
Building on this input structure, we employ a five-layer Graph Neural Network architecture tailored for task fingerprinting from fMRI data. The network starts with a GATv2Conv layer that re-weights each node’s neighbours, emphasising task-relevant connections [54] (input $200$, out $32$ features with $4$ attention heads, dropout $0.30$). A parallel linear skip path (output $128$ channels) is summed with the convolution result so the original signal remains visible to deeper layers, improving gradient flow and protecting against over-smoothing [55]. TopKPooling then selectively retains the most informative nodes (ratio $0.8$), reducing graph complexity and focusing computation on salient sub-structures. This sparsification step improves both efficiency and performance. A second GATv2Conv block, again accompanied by a skip connection, further refines the retained features and maps them into a more discriminative space for task recognition (input $128$, out $32$ features with $2$ heads, dropout $0.30$; $32\times 2=64$ output channels; skip output $64$ channels). Global-mean pooling aggregates these refined node features across the graph, producing a single $64$-dimensional vector. A linear head of size $64\times 8$, followed by log-softmax, converts this vector into task probabilities for training and evaluation. During training, we randomly drop $10\%$ of edges to enhance robustness. The overall model structure is visualised in Figure 3. The model is trained using supervised learning with data batches drawn from $\widehat{\mathcal{R}}$. The rows of $R^{D}_{s,h}$ serve as the input features for each node, while $h$ is used as the corresponding fMRI task label. Detailed training parameters are provided in the Appendix.
@@ 4 Results
Method CM+MD&P CM+CoR CM+FN CM+GNN FC+CoR FC+MD&P FC+FN FC+GNN TSF PH-H0
Rest1 LR 95.396% 58.397% 42.251% $5\%$ 45.439% 69.736% 33.952% $5\%$ 88.7% 85.6%
Rest2 LR 96.334% 61.295% 47.175% $5\%$ 45.436% 70.929% 37.574% $5\%$ 89.4% 86.2%
Rest1 RL 94.970% 57.118% 41.333% $5\%$ 41.603% 67.945% 38.698% $5\%$ 87.9% 84.7%
Rest2 RL 94.203% 58.978% 43.463% $5\%$ 46.803% 67.775% 36.737% $5\%$ 88.5% 84.3%
@@
PH-H0: Persistent Homology(H0 Landscape) [47] TSF: Tangent-Space Fingerprinting [46]
@@ 4.1 Dataset and Modeling
Specifically, we use the minimally pre-processed resting-state runs released by the UCSD Library [59, 60] based on the HCP Young Adult raw data [61]. These scans have already undergone the standard HCP pipeline [62], which includes gradient-distortion and motion correction, EPI distortion correction, alignment to MNI space, high-pass filtering with a 2000-s cut-off, and ICA-FIX denoising [63]. After parcellation with the Schaefer-100 atlas, each run yields a $100\times 1190$ matrix of regional time-series sampled every 0.720.72[64].
Figure 6: Left: Subject Fingerprinting based on CM+MD&P: The identification accuracy is compared with (black line) using the FC+CoR method. Right: Task Fingerprinting based on CM+GNN: The figure shows the correct/incorrect classification for the same/different fMRI tasks. The y-axis is split. The top segment (95–100%) shows the near-perfect accuracies, while the bottom segment (0–2%) zooms in on the small error rates.